The home states of a certain group of people are distributed as follows: 53 percent are from MISSOURI, 27 percent are from KANSAS, and 20 percent are from IOWA. (No one in the group had a home state other than one of these three.) (Note: Your answer to the question below should be rounded to three decimal places.) Suppose we randomly select a person from this group. What is the expected value of the number of letters in the selected person's home state?
Answer #1

53 percent are from MISSOURI, 27 percent are from KANSAS, and 20 percent are from IOWA.

X: Number of letters

x
8
6
4

We have to find the expected value of the number of letters in the selected person's home state.

Formula is

E(X)=\sum x*p

x p xp
8 0.53 4.24
6 0.27 1.62
4 0.2 0.8
Total 6.66

E(X)=\sum x*p=6.66

The expected value of the number of letters in the selected person's home state is 6.66